What is the % of Na in a 0.721g detergent sample. Final yield is 0.320g NaP2O4.


1
Expert's answer
2021-02-09T04:36:36-0500

First, calculate the mass of sodium in the NaP2O4 sample:

m(Na) = m(NaP2O4) × Mr(Na) / Mr(NaP2O4) = 0.320 g × 23 g / 149 g/mol = 0.05 g

Second, calculate % of Na in a detergent sample:

% Na = (m(Na) / Msample) × 100 % = (0.05 g / 0.721 g) × 100 % = 3.6 %


Answer: 3.6 %